4. Calculate the percentage of the Real GDP components for 2016 in the table format below. (Show calculations) (Show calculations) (4 points) Billions of chained Percentage of Real GDP Components (2009) dollars 16,716.2 Real GDP Consumption 11,572.1 Investment 2,858.3 Government Expenditure 2,900.2 Net Export - 586.3
Answer #1

Percentage of Real GDP Components :

1) Consumption - (11572.1 ÷ 16716.2 ) × 100 = 69.3%

2) Investment - ( 2858.3 ÷ 16716.2 ) × 100 = 17.2%

3) Government Expenditure - ( 2900.2 ÷ 16716.2 ) × 100 = 17.3%

4) Net Exports - ( -568.3 ÷ 16716.2 ) × 100 = - 3.4%

Best of Luck !! !!